# Degrees of the Wasserstein Distance to Small Toric Models

### v1

@@ 3.1. Computational experiments
$a,b$
$f$-vector
Codimension 3
Codimension 2
Codimension 1
1,2
(8, 24, 32, 16)
$\begin{cases}1:8\\ 2:8\\ 3:8\end{cases}$
$\begin{cases}1:4\\ 2:4\\ 3:12\\ 4:12\end{cases}$
$\begin{cases}3:16\end{cases}$
1,3
(10, 40, 80, 80, 32)
$\begin{cases}1:16\\ 2:24\\ 3:24\\ 4:16\end{cases}$
$\begin{cases}0:2\\ 2:2\\ 3:16\\ 4:46\\ 6:14\end{cases}$
$\begin{cases}3:32\end{cases}$
1,4
(12, 60, 160, 240, 192, 64)
$\begin{cases}1:32\\ 2:48\\ 3:32\\ 4:64\\ 5:64\end{cases}$
$\begin{cases}5:80\\ 6:16\\ 7:48\\ 8:48\\ \end{cases}$
$\begin{cases}5:64\end{cases}$
1,5
(14, 84, 280, 560, 672, 448, 128)
$\begin{cases}1:64\\ 2:96\\ 3:96\\ 4:224\\ 5:128\\ 6:64\end{cases}$
$\begin{cases}3:12\\ 5:64\\ 6:64\\ 7:52\\ 8:192\\ 10:64\\ \end{cases}$
$\begin{cases}5:128\end{cases}$
2,3
(12, 60, 160, 240, 192, 64)
$\begin{cases}1:32\\ 2:80\\ 3:32\\ 4:32\\ 5:64\end{cases}$
$\begin{cases}4:16\\ 5:16\\ 6:48\\ 7:80\\ 8:32\\ \end{cases}$
$\begin{cases}5:64\end{cases}$
2,4
(14, 84, 280, 560, 672, 448, 128)
$\begin{cases}1:64\\ 2:160\\ 3:64\\ 4:128\\ 5:64\\ 6:192\end{cases}$
$\begin{cases}2:4\\ 4:4\\ 6:44\\ 7:128\\ 8:96\\ 9:64\end{cases}$
$\begin{cases}6:128\end{cases}$
@@
Acknowledgements
We thank Bernd Sturmfels and Jose Israel Rodriguez for helpful conversations.

### v2

@@ Acknowledgements
$a,b$ $f$-vector Codimension 3 Codimension 2 Codimension 1
1,2 (8, 24, 32, 16) $\begin{cases}1:8\\ 2:8\\ 3:8\end{cases}$$\begin{cases}1:4\\ 2:4\\ 3:12\\ 4:12\end{cases}$$\begin{cases}3:16\end{cases}$
1,3 (10, 40, 80, 80, 32) $\begin{cases}1:16\\ 2:24\\ 3:24\\ 4:16\end{cases}$$\begin{cases}0:2\\ 2:2\\ 3:16\\ 4:46\\ 6:14\end{cases}$$\begin{cases}3:32\end{cases}$
1,4 (12, 60, 160, 240, 192, 64) $\begin{cases}1:32\\ 2:48\\ 3:32\\ 4:64\\ 5:64\end{cases}$$\begin{cases}5:80\\ 6:16\\ 7:48\\ 8:48\\ \end{cases}$$\begin{cases}5:64\end{cases}$
1,5 (14, 84, 280, 560, 672, 448, 128) $\begin{cases}1:64\\ 2:96\\ 3:96\\ 4:224\\ 5:128\\ 6:64\end{cases}$$\begin{cases}3:12\\ 5:64\\ 6:64\\ 7:52\\ 8:192\\ 10:64\\ \end{cases}$$\begin{cases}5:128\end{cases}$
2,3 (12, 60, 160, 240, 192, 64) $\begin{cases}1:32\\ 2:80\\ 3:32\\ 4:32\\ 5:64\end{cases}$$\begin{cases}4:16\\ 5:16\\ 6:48\\ 7:80\\ 8:32\\ \end{cases}$$\begin{cases}5:64\end{cases}$
2,4 (14, 84, 280, 560, 672, 448, 128) $\begin{cases}1:64\\ 2:160\\ 3:64\\ 4:128\\ 5:64\\ 6:192\end{cases}$$\begin{cases}2:4\\ 4:4\\ 6:44\\ 7:128\\ 8:96\\ 9:64\end{cases}$$\begin{cases}6:128\end{cases}$